OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 128 and a3 = 8?

OpenStudy (anonymous):

i just need a formula, is all

OpenStudy (kamille):

so, whats the problem? Use a formula?

OpenStudy (anonymous):

yes, that's what i just asked for..

OpenStudy (anonymous):

@nincompoop

OpenStudy (kamille):

oh, so you can google it, can't you? http://en.wikipedia.org/wiki/Geometric_progression#Elementary_properties

OpenStudy (anonymous):

i didnt understand it which is why I'm here....seeking help

OpenStudy (kamille):

So, you know what formula is \[a ^{n}=a _{1}q ^{n-1}\]

OpenStudy (anonymous):

@rajathsbhat @Kamille i don't get the formula is my problem

OpenStudy (anonymous):

does this help ? http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (anonymous):

the common ratio, how do you find that?

OpenStudy (kamille):

oh, it should be \[a _{n}=a _{1}q^{n-1}\] Well, to get the formula is very easy. Let's say, that there is a number a1 then a2=a1*q, a3=a2*q etc, do you understand why?

OpenStudy (anonymous):

\[\Large a_{3}=a_{1}q^{2}\] right? you have a3 and a1. Substitute them into the equation to find q. You can then find a6.

OpenStudy (kamille):

geometrical progression means that very is a sequance, where a1;a2(a1*q);a3(a2*q), etc when q=a2/a1

OpenStudy (anonymous):

=0.125

OpenStudy (anonymous):

isn't it 0.25 or 1/4?

OpenStudy (kamille):

it is 1/4

OpenStudy (kamille):

well, why it can't be -1/4?

OpenStudy (kamille):

q must be >0?

OpenStudy (anonymous):

yep,equally possible

OpenStudy (anonymous):

no its 0.125

OpenStudy (anonymous):

thanks guys

OpenStudy (kamille):

you have forgoten to square root the answer, yummydum

OpenStudy (anonymous):

i put 0.125 and it was right...